OpenStudy (anonymous):

converting standard form to slope-intercept form help i know how to do it, but I am having troubles with this one 2x + 3y = 1,200

OpenStudy (anonymous):

i know we start by subtracting 2x on both sides of the equation to get 3y = 1,200 - 2x

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

or do we add -2x to get 3y = 1,200 + -2x

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

see what i got is y = 400 + -2/3x???? is that right

OpenStudy (anonymous):

or y = 400 - 2/3x but with the formula y = mx + b, should it be the addition sign

OpenStudy (sleepyjess):

It should be y = 400 - 2/3x, but switch around 400 and -2/3x to make y = -2/3x + 400

OpenStudy (anonymous):

oh I see, so I was right?

OpenStudy (sleepyjess):

yep

OpenStudy (anonymous):

So the slope = -2/3 and the y-intercept is 400?

OpenStudy (anonymous):

i am in no rush :) get here when you want to help

OpenStudy (sleepyjess):

Oh, sorry, I was trying to get help on a question of my own haha yes, that is correct :)

OpenStudy (anonymous):

that is why i said i am in no rush, i saw ;p and thanks for your help

OpenStudy (sleepyjess):

no problem :)

OpenStudy (anonymous):

and in function notation form it's f(x) = -2/3x + 400, correct @sleepyjess

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

oh so that is how you delete a reply ;p haha i didn't know that

OpenStudy (anonymous):

ok! thans for your help

OpenStudy (sleepyjess):

:)